Find the future value of a five-year $113,000 investment that pays 10.00 percent and that has the following compounding periods: (Do not round intermediate calculations, round final answers to 2 decimal places, e.g. 15.25.) a. Quarterly b. Monthly c. Daily d. Continuous
Answer #1

Future value = Present value x (1+i=n)^nt Future value = FV = = $113,000 X (1+10%/4)^(4*5) $185,163.66 Future value = FV = $1